# AI-Powered Episodic Future Thinking

### v1

@@ Abstract
Episodic Future Thinking (EFT) is an intervention that involves vividly imagining personal future events and experiences in detail. It has shown promise as an intervention to reduce delay discounting—the tendency to devalue delayed rewards in favor of immediate gratification—and to promote behavior change in a range of maladaptive health behaviors. We present EFTeacher, an AI chatbot powered by the GPT-4-Turbo large language model, designed to generate EFT cues for users with lifestyle-related conditions. To evaluate the chatbot, we conducted a user study that included usability assessments and user evaluations based on content characteristics questionnaires, followed by semi-structured interviews. The study provides qualitative insights into participants’ experiences and interactions with the chatbot and its usability. Our findings highlight the potential application of AI chatbots based on Large Language Models (LLMs) in EFT interventions, and offer design guidelines for future behavior-oriented applications.
@@ 1 Introduction
Approximately 40% of adults in the United States are currently living with obesity, with projections suggesting that this figure will reach 50% by 2030 [1, 2]. Although some progress has been made in addressing obesity, many communities still lack access to prevention and treatment resources, worsening health disparities related to obesity [3]. Losing weight successfully depends on making choices today that are influenced by long-term goals, as most of the health benefits from weight loss come after months or even years of maintaining consistent healthy habits.
@@
Episodic future thinking (EFT) involves vividly imagining personal future events and experiences in detail [4]. EFT has been increasingly used as an intervention for behavior change in various maladaptive health behaviors and populations with lifestyle-related conditions to reduce delay discounting [5]. Delay discounting (DD) refers to the tendency to choose immediate reward over greater, future benefits [6]. EFT interventions usually involve participants identifying several positive future events they look forward to and plan to attend at different times. They write or describe these events in detail and are encouraged to vividly imagine them when making decisions involving future trade-offs [7]. Personalized positive future cues have been shown to significantly impact decision-making by encouraging individuals to consider longer-term consequences. These cues help expand the "temporal window," which refers to the ability to think further into the future when evaluating choices [8, 9, 10].
@@
Research indicates that obesity is linked to high levels of DD. For instance, high levels of DD are both currently and over time linked to having a higher body mass index (BMI), consuming more calories, and engaging in less physical activity [11, 12, 13]. Additionally, experimental studies demonstrate that interventions aimed at altering DD lead to related improvements in dietary habits, further supporting the idea of a functional connection between the two [13, 14, 15, 16].
@@
Laboratory studies have shown that engaging in EFT for a short time can effectively reduce DD in adults with obesity [17, 18, 19, 20, 21], and a range of other clinical disorders [22] . This means EFT helps people prioritize long-term goals over immediate rewards. Studies also demonstrate that brief EFT sessions can decrease the purchase and consumption of unhealthy, high-calorie foods both in controlled lab settings and in real-life situations, such as grocery shopping or dining in public [14, 23, 16, 24].
Building on these findings, EFT has been adapted into a remotely delivered intervention as part of clinical trials for managing type 2 diabetes. This intervention aims to help individuals adhere to lifestyle changes recommended for treating obesity and its comorbidities [25, 26, 5]. EFT works by helping patients stay focused on their long-term health goals while reducing the appeal of immediate temptations like unhealthy foods or inactivity, which could derail their progress.
@@
There are two common approaches to creating EFT cues for studies on delay discounting and health behaviors: interview-guided and survey-guided methods [17, 27]. Both methods share some similarities in how the cues are generated. For instance, researchers may give specific instructions based on the study’s goals and participants’ behavior [5]. In clinical practice, having clinicians conduct interviews to guide patients through EFT is costly, requires significant training to ensure treatment fidelity, and limits when and where patients can access the intervention. While self-administered, survey-based approaches address these challenges by reducing costs and increasing accessibility, they introduce new concerns about maintaining consistency and effectiveness in treatment (e.g., due to terse or off-topic cues). To address these issues, this study aims to evaluate an artificial intelligence (AI)-based method for delivering EFT. This approach would improve consistency, scalability, accessibility, and effectiveness, allowing EFT to be implemented on a much larger scale while minimizing costs and other barriers. We aim to enhance and evaluate an existing AI-based chatbot [28], referred to as EFTeacher, for EFT intervention. EFTeacher is a chatbot built on GPT-4-turbo, a Large Language Model (LLM). EFTeacher is designed to assist participants in identifying personally meaningful future events, and in creating for them detailed text descriptions, or cues, for those events.
@@
  • RQ1: What are users’ overall experiences of using EFTeacher?
  • RQ2: What are users impressions of using EFTeacher?
@@
To address RQ1 and RQ2, we adopt a quantitative approach utilizing the System Usability Scale (SUS) [29], and an analysis of specific content characteristics [30]. We answer RQ3 by using thematic analysis [31, 32].
@@ 2 Related Work
Conversational agents have emerged as valuable tools for delivering cost-effective and practical behavioral interventions. Studies demonstrate their effectiveness in promoting physical activity, encouraging healthy diets [33, 34], and supporting weight loss efforts [35]. The design of engaging and impactful chatbot interventions remains a crucial focus for maximizing their potential [36]. Furthermore, AI-driven conversational agents have been used in behavior change systems to create more interactive and engaging interventions [37, 38, 39, 40].
@@
Human-computer interaction (HCI) technologies have shown promise in fostering mindfulness and enhancing well-being [41]. Studies highlight its ability to increase relaxation, focus, and self-awareness, thereby improving decision-making and motivating behavior change [41]. Mindful interaction with technology emphasizes the importance of reflection and deliberate choices during digital interactions, further enhancing personal growth and well-being. As a result, researchers have explored various approaches to designing conversational agents tailored to specific behavioral goals.
@@
Physical Activity and Healthy Eating: Olafsson et al. [42] compared the impact of humorous versus non-humorous virtual agents in promoting exercise and healthy eating, finding humor to enhance motivation and engagement significantly. Similarly, Sun et al. [43] investigated the impact of humor in chatbot-delivered interventions designed to increase physical activity. The study focused on user interactions with a chatbot, analyzing levels of participation, motivation, and physical activity. The findings showed that the use of humor in the chatbot helped users become more physically active. Paay et al. [44] highlighted the role of digital personal assistants (DPAs) in motivating physical activity such as exercise, with virtual rewards being particularly effective. Kocielnik et al. [45] introduced a reflection tool called the “Reflection Companion,” which encourages users to notice patterns in their physical activity and plan future actions. The system uses short daily conversations and visual data representations to help users recognize their habits and strategize for improvement, underscoring the value of reflective practices in promoting healthy behaviors.
@@
Smoking Cessation: Alphonse et al. [46] revealed that human-like traits in chatbots promote accountability and engagement, making them effective tools for smoking cessation. Their qualitative study explored smokers’ experiences with Quit Coach, a quick-response chatbot integrated into a smoking cessation app. The findings suggest that the chatbot’s design increased users’ sense of accountability and engagement, supporting the use of chatbots in smoking cessation interventions. Kjaerulff et al. [47] explored voice-based user interfaces (VUIs) that facilitated mindfulness techniques for smoking cessation. Participants used the VUI to track cravings, engage in mindfulness exercises, or reflect on smoking habits by listening to pre-recorded audio responses. Results indicated that both VUIs and augmented VUIS (AVUIs) show promise in supporting smoking cessation and proposed integrating Large Language Models (LLMs) for enhanced personalization and adaptability.
@@
Applications in Chronic Disease Management: Conversational agents have also shown promise in managing chronic conditions like diabetes. Bin et al. [48] proposed a chatbot design framework that addresses limitations in existing systems. This framework was developed using surveys, interviews, and a review of current healthcare chatbots. By incorporating insights from the interviews, the framework was designed to meet user needs and preferences, thereby improving the chatbot’s effectiveness in preventing type 2 diabetes. Baptista [49] evaluated the acceptability of an embodied conversational agent (ECA) for type 2 diabetes self-management. Using both survey data and participant interviews, the study found that users appreciated the agent’s friendly and motivational support, though improvements in communication and appearance were recommended, highlighting the potential of ECAs in chronic disease management.
@@
Supporting Well-Being and Mindfulness in Food-Related Activities: Studies have also examined how conversational agents can enhance well-being during food-related activities. Parra et al. [50] explored the potential of agents to promote mindfulness during meals, especially following the impact of COVID-19 on eating habits. The study began by examining how the COVID-19 pandemic affected young adults’ eating habits, noting its negative impact on well-being. Based on these findings, the researchers developed two interactive systems. Both systems demonstrated that conversational agents could make these activities more enjoyable and foster mindfulness. The researchers aim to develop more advanced agents using LLMs and activity recognition to adapt to users’ needs effectively. Zhang et al. [51] presented a pilot study investigating how voice assistants can encourage mindful eating. Researchers created two versions of a voice assistant app: one with a friendly approach and another with a counseling approach. They tested these apps to evaluate user preferences and how the assistant’s social role influenced their mindful eating experience. The study found that participants preferred the "friend" version, offering insights for designing voice assistants that support mindfulness practices.
@@
Advancing Conversational Agents for Personal Growth: Several studies focus on refining conversational agents to support broader personal development goals. Saravanan et al. [52] explored the design and evaluation of a conversational system for a social robot employing motivational interviewing to encourage healthy eating. Their findings demonstrated that referencing past interactions in motivational interviewing sessions improved user engagement and behavior change. Participants interacted with the virtual robot over three sessions, during which their motivation levels and goal progress were measured. Results indicated that varied references to previous sessions significantly increased motivation for behavior change. Meyer et al. [53] highlighted the potential of conversational agents trained to recognize behavior-change language using motivational interview concepts. Browne et al. [54] developed and tested a conversational AI health coaching system integrated into an app. The study evaluated the system’s performance, user interaction, and the effectiveness of prompt engineering techniques. They assessed the app’s impact on mental well-being, finding improvements in positive affect, reduced negative affect, and enhanced self-reflection. The study concludes with a discussion of the app’s usability, user feedback, and suggestions for future improvements.
@@
Large Language Models in Conversational Systems: Recent advancements in LLMs have further expanded the capabilities of conversational agents. Arakawa et al. [55] explored LLM-powered chatbots in executive coaching, emphasizing their accessibility and logical reasoning while highlighting the importance of human-AI collaboration to promote deep self-reflection. James et al. [56] evaluated LLMs for generating personalized goals in gamified mHealth apps. A study comparing LLM-generated and pre-determined goals highlighted the scalability and efficiency of LLM-driven content creation. Nepal et al. [57] introduced the MindScape app, which uses AI-generated journaling prompts tailored to smartphone behavioral data. The study showed improvements in mental well-being and self-reflection, emphasizing the app’s potential for enhancing traditional journaling for effective interventions that promote self-awareness and holistic well-being.
Our study aims to perform a qualitative evaluation of EFTeacher, which is based on a Large Language Model (LLM) for EFT cue generation. To the best of our knowledge, we are the first group to leverage LLMs for EFT cue generation, which has potential applications in interventions addressing maladaptive health behaviors.
3 Study Method
@@
We developed an AI chatbot [28], EFTeacher, which is modeled after the clinician-led EFT interviews [17, 16, 20] and self-administered surveys [19, 18, 58] used previously. We implemented a website [59] for a single-session study in which participants were given a username and password and could log into the website as shown in Figure 1 to complete the AI-assisted EFT interview. We used GPT-4-Turbo [60] through Microsoft Azure, which is a university-approved cloud service for EFTteacher. As shown in Figure 1, memory, in the form of a list, is the history of the conversation which is passed into the prompt along with the user query (message) so the chatbot can generate an output. As depicted in the diagram, the memory stores both the responses generated by the chatbot and the corresponding user queries. Each interaction between the human and GPT-4 is added to this list so that it can be used in the next input for the model.
@@
On the chat page of the website, depicted in Figure LABEL:wireframe, users interacted with the chatbot, which guided them in generating a set of EFT cue texts. EFTeacher first introduced EFT cues to participants. They were then provided with clear criteria for crafting a good EFT cue. Participants were guided to ensure their cues met specific criteria: the event should be something they look forward to, vividly imaginable, confined to a single 24-hour period, and written in the present tense as if they are currently experiencing it. To help them understand, participants were shown examples of well-crafted EFT cues that met all the criteria, as well as poor examples. They were then prompted to create their own EFT cues for specific timeframes, starting from one month into the future, ensuring their events were realistic, specific, and followed the given guidelines. This step-by-step approach ensured participants understood EFT cues and could create effective ones.
@@
For each time-frame, participants first provided brief contextual statements about the event. Then they answered a series of standardized questions to add specific details, such as where the event was taking place, who was present, what actions were occurring, and how they felt in that moment. EFTeacher then combined these brief statements and detailed responses into a coherent cue, typically 3-4 sentences long. If participants gave incomplete or unrelated answers, EFTeacher prompted them to revise their responses to ensure the quality and relevance of the cues. Once an EFT cue was generated, EFTeacher sought feedback from the participant to adapt the cue based on their preferences, ensuring it aligned with their personal perspective and needs. The user and chatbot interacted iteratively until a desired EFT cue was generated and accepted by the user. Once finalized, the EFT cue was extracted through a predefined routine and displayed in a dedicated box (cue box), as shown in Figure LABEL:wireframe. As the conversation progressed and an EFT cue was generated for any time-frame, users could view the cues in the cue box, navigate back to previous cues, and review them. This allowed users to keep track of which events had already been discussed during the EFT generation For this study, once users completed covering events for seven different time-frames (one month, three months, one year, three years, five years, and ten years), the conversation concluded.
@@
Upon completing the conversation, the chatbot’s logs, including the entire conversation (referred to as memory), were saved as a file in the database. A designated routine stored the extracted EFT cues in the database, and then sent them to Qualtrics. Following this, users were redirected to the next page, where they evaluated the system’s usability using the System Usability Scale (SUS) [29] questions. After completing these questions, they were redirected to the Qualtrics website. In Qualtrics, participants were presented with the generated EFT cue texts by EFTeacher, organized by the different time-frames. For each cue, participants were asked to evaluate its characteristics on a 5-point Likert scale, ranging from 1 (not at all) to 5 (extremely). They rated each cue based on four key aspects: Liking, Importance, Excitement, and Vividness. Afterward, the participants were presented with a survey to provide additional feedback about their experience by answering open-ended questions, as explained in detail in Section 3.2.

### v2

# template.bib

@@ Abstract
Episodic Future Thinking (EFT) involves vividly imagining personal future events and experiences in detail. It has shown promise as an intervention to reduce delay discounting—the tendency to devalue delayed rewards in favor of immediate gratification—and to promote behavior change in a range of maladaptive health behaviors. We present EFTeacher, an AI chatbot powered by the GPT-4-Turbo large language model, designed to generate EFT cues for users with lifestyle-related conditions. To evaluate the feasibility and usability of EFTeacher, we conducted a mixed-methods study that included usability assessments, user evaluations based on content characteristics questionnaires, and semi-structured interviews. Qualitative findings indicate that participants perceived EFTeacher as communicative and supportive through an engaging dialogue. The chatbot facilitated imaginative thinking and reflection on future goals. Participants appreciated its adaptability and personalization features, though some noted challenges such as repetitive dialogue and verbose responses. Our findings underscore the potential of large language model-based chatbots in EFT interventions targeting maladaptive health behaviors.
@@ 1 Introduction
Approximately 40% of adults in the United States are currently living with obesity, with projections suggesting that this figure will reach 50% by 2030 [lieberman2022state, revels2017predicting]. Although some progress has been made in addressing obesity, many communities still lack access to prevention and treatment resources, worsening health disparities related to obesity [washington2023disparities]. Losing weight successfully depends on making choices today that are influenced by long-term goals, as most of the health benefits from weight loss come after months or even years of maintaining consistent healthy habits.
@@
Episodic future thinking (EFT) involves vividly imagining personal future events and experiences in detail [atance2001episodic]. EFT has been increasingly used as an intervention for behavior change regarding various maladaptive health behaviors, and for populations with lifestyle-related conditions, to reduce delay discounting [brown2022putting]. Delay discounting (DD) refers to the tendency to choose immediate reward over greater, future benefits [bickel2012excessive]. EFT interventions usually involve participants identifying several positive future events they look forward to and plan to attend at different times. They write or describe these events in detail and are encouraged to vividly imagine them when making decisions involving future trade-offs [bickel2017toward]. Personalized positive future cues have been shown to significantly impact decision-making by encouraging individuals to consider longer-term consequences. These cues help expand the “temporal window,” which refers to the ability to think further into the future when evaluating choices [bulley2017influence, chiou2017episodic, kaplan2016effects].
@@
Research indicates that obesity is linked to high levels of DD. For instance, high levels of DD are both currently and over time linked to having a higher body mass index (BMI), consuming more calories, and engaging in less physical activity [bickel2021temporal, appelhans2019delay, garza2016impulsivity]. Additionally, experimental studies demonstrate that interventions aimed at altering DD lead to related improvements in dietary habits, further supporting the idea of a functional connection between the two [garza2016impulsivity, hollis2019episodic, snider2020reinforcer, hollis2020ecological].
@@
Laboratory studies have shown that engaging in EFT for a short time can effectively reduce DD in adults with obesity [daniel2013future, brown2023episodic, stein2021bleak, bickel2020does, stein2017think], and a range of other clinical disorders [rosch2022promoting] . This means EFT helps people prioritize long-term goals over immediate rewards. Studies also demonstrate that brief EFT sessions can decrease the purchase and consumption of unhealthy, high-calorie foods both in controlled lab settings and in real-life situations, such as grocery shopping or dining in public [hollis2019episodic, hollis2020mothers, hollis2020ecological, o2016episodic]. Building on these findings, EFT has been adapted into a remotely delivered intervention as part of clinical trials for managing type 2 diabetes. This intervention aims to help individuals adhere to lifestyle changes recommended for treating obesity and its comorbidities [sze2015web, epstein2022effects, brown2022putting]. EFT works by helping patients stay focused on their long-term health goals while reducing the appeal of immediate temptations like unhealthy foods or inactivity, which could derail their progress.
@@
There are two common approaches to creating EFT cues for studies on delay discounting and health behaviors: interview-guided and survey-guided methods [daniel2013future, sze2017bleak]. Both methods share some similarities in how the cues are generated. For instance, researchers may give specific instructions based on the study’s goals and participants’ behavior [brown2022putting]. In clinical practice, having clinicians conduct interviews to guide patients through EFT is costly, requires significant training to ensure treatment fidelity, and limits when and where patients can access the intervention. While self-administered, survey-based approaches address these challenges by reducing costs and increasing accessibility, they introduce new concerns about maintaining consistency and effectiveness in treatment (e.g., due to terse or off-topic cues). Although the survey method was designed as a more scalable alternative to the human interview, poor quality cues in this self-administered method (e.g., vague or non-episodic) may reduce treatment fidelity and thus efficacy of the intervention. Although EFT is a promising behavioral intervention to facilitate weight loss and dietary improvement [daniel2013future, brown2023episodic, stein2021bleak, bickel2020does, stein2017think], significant effects on health behaviors are not always observed [bickel2020does, stein2021bleak]. Thus, research is needed to examine whether the method used to generate cues moderates these behavioral effects and user perceptions.
To address these issues, this study aims to evaluate an artificial intelligence (AI)-based method for delivering EFT. This approach would improve consistency, scalability, accessibility, and effectiveness, allowing EFT to be implemented on a much larger scale while minimizing costs and other barriers. We aim to enhance and evaluate an existing prototype AI-based chatbot [ahmadi2024ai], referred to as EFTeacher, for EFT intervention. EFTeacher is a chatbot built on GPT-4-turbo, a Large Language Model (LLM). EFTeacher is designed to assist participants in identifying personally meaningful future events, and in creating for them detailed text descriptions, or cues, for those events.
@@
  • RQ1: How do users evaluate the usability of EFTeacher?
  • RQ2: What are users’ impressions of the EFT cue texts generated by EFTeacher?
@@
To address RQ1 and RQ2, we used the System Usability Scale (SUS) [brooke1996sus], recognizing that usability is a necessary foundation for any tool to be effective, including AI systems for EFT interventions. For RQ2, we analyzed specific content characteristics of the generated EFT cues generated by EFTeacher [rung2020translating], as these measures have been used in previous studies [rung2020translating] to assess the effects of EFT cue texts. We address RQ3 using thematic analysis [braun2006using, bowman2023using], a widely used qualitative method for identifying and analyzing patterns in user perceptions and experiences.
To the best of our knowledge, this is the first study to apply LLMs for EFT cue generation—a key procedural component in behavioral interventions targeting delay discounting. Prior work has not developed, evaluated, or validated an AI-based method for eliciting EFT cues. This proof-of-concept study aims to evaluate the usability and acceptability of this novel approach, in line with prior work emphasizing the importance of user-centered and theory-based approaches to effectively develop AI-guided systems [nadarzynski2019acceptability].
@@ 2 Related Work
Conversational agents have emerged as valuable tools for delivering cost-effective and practical behavioral interventions. Studies demonstrate their effectiveness in promoting physical activity, encouraging healthy diets [zhang2020artificial, oh2021systematic], and supporting weight loss efforts [chew2022use]. The design of engaging and impactful chatbot interventions remains a crucial focus for maximizing their potential [cole2019understanding]. Furthermore, AI-driven conversational agents have been used in behavior change systems to create more interactive and engaging interventions [li2023exploring, kumar2023exploring, jo2023understanding, nepal2024contextual].
@@
Human-computer interaction (HCI) technologies have shown promise in fostering mindfulness and enhancing well-being [terzimehic2019review]. Studies highlight their ability to increase relaxation, focus, and self-awareness, thereby improving decision-making and motivating behavior change [terzimehic2019review]. Mindful interaction with technology emphasizes the importance of reflection and deliberate choices during digital interactions, further enhancing personal growth and well-being. As a result, researchers have explored various approaches to designing conversational agents tailored to specific behavioral goals.
@@
Physical Activity and Healthy Eating: Olafsson et al. [olafsson2020motivating] compared the impact of humorous versus non-humorous virtual agents in promoting exercise and healthy eating, finding humor can enhance motivation and engagement significantly. Similarly, Sun et al. [sun2024can] investigated the impact of humor in chatbot-delivered interventions designed to increase physical activity. The study focused on user interactions with a chatbot, analyzing levels of participation, motivation, and physical activity. The findings showed that the use of humor in the chatbot helped users become more physically active. Paay et al. [paay2022can] highlighted the role of digital personal assistants (DPAs) in motivating physical activity such as exercise, with virtual rewards being particularly effective.
Kocielnik et al. [kocielnik2018reflection] introduced a reflection tool called the “Reflection Companion,” which encourages users to notice patterns in their physical activity and plan future actions. The system uses short daily conversations and visual data representations to help users recognize their habits and strategize for improvement, underscoring the value of reflective practices in promoting healthy behaviors.
@@
Smoking Cessation: Alphonse et al. [alphonse2022exploring] revealed that human-like traits in chatbots promote accountability and engagement, making them effective tools for smoking cessation. Their qualitative study explored smokers’ experiences with Quit Coach, a quick-response chatbot integrated into a smoking cessation app. The findings suggest that the chatbot’s design increased users’ sense of accountability and engagement, supporting the use of chatbots in smoking cessation interventions. Kjaerulff et al. [kjaerulff2024exploring] explored voice-based user interfaces (VUIs) that facilitated mindfulness techniques for smoking cessation. Participants used the VUI to track cravings, engage in mindfulness exercises, or reflect on smoking habits by listening to pre-recorded audio responses. Results indicated that both VUIs and augmented VUIS (AVUIs) show promise in supporting smoking cessation and proposed integrating Large Language Models (LLMs) for enhanced personalization and adaptability.
@@
Applications in Chronic Disease Management: Conversational agents have also shown promise in managing chronic conditions like diabetes. Bin et al. [bin2023towards] proposed a chatbot design framework that addresses limitations in existing systems. This framework was developed using surveys, interviews, and a review of current healthcare chatbots. By incorporating insights from the interviews, the framework was designed to meet user needs and preferences, thereby improving the chatbot’s effectiveness in preventing type 2 diabetes. Baptista [baptista2020acceptability] evaluated the acceptability of an embodied conversational agent (ECA) for type 2 diabetes self-management. Using both survey data and participant interviews, the study found that users appreciated the agent’s friendly and motivational support, though improvements in communication and appearance were recommended, highlighting the potential of ECAs in chronic disease management.
@@
Supporting Well-Being and Mindfulness in Food-Related Activities: Studies have also examined how conversational agents can enhance well-being during food-related activities. Parra et al. [parra2023enhancing] explored the potential of agents to promote mindfulness during meals, especially following the impact of COVID-19 on eating habits. The study began by examining how the COVID-19 pandemic affected young adults’ eating habits, noting its negative impact on well-being. Based on these findings, the researchers developed two interactive systems. Both systems demonstrated that conversational agents could make these activities more enjoyable and foster mindfulness. The researchers aim to develop more advanced agents using LLMs and activity recognition to adapt to users’ needs effectively. Zhang et al. [zhang2023facilitating] presented a pilot study investigating how voice assistants can encourage mindful eating. Researchers created two versions of a voice assistant app: one with a friendly approach and another with a counseling approach. They tested these apps to evaluate user preferences and how the assistant’s social role influenced their mindful eating experience. The study found that participants preferred the “friend” version, offering insights for designing voice assistants that support mindfulness practices.
@@
Advancing Conversational Agents for Personal Growth: Several studies focus on refining conversational agents to support broader personal development goals. Saravanan et al. [saravanan2022giving] explored the design and evaluation of a conversational system for a social robot employing motivational interviewing to encourage healthy eating. Their findings demonstrated that referencing past interactions in motivational interviewing sessions improved user engagement and behavior change. Participants interacted with the virtual robot over three sessions, during which their motivation levels and goal progress were measured. Results indicated that varied references to previous sessions significantly increased motivation for behavior change. Meyer et al. [meyer2023towards] highlighted the potential of conversational agents trained to recognize behavior-change language using motivational interview concepts. Browne et al. [browne2024reflective] developed and tested a conversational AI health coaching system integrated into an app. The study evaluated the system’s performance, user interaction, and the effectiveness of prompt engineering techniques. They assessed the app’s impact on mental well-being, finding improvements in positive affect, reduced negative affect, and enhanced self-reflection. The study concludes with a discussion of the app’s usability, user feedback, and suggestions for future improvements.
@@
Large Language Models in Conversational Systems: Recent advancements in LLMs have further expanded the capabilities of conversational agents. Arakawa et al. [arakawa2024coaching] explored LLM-powered chatbots in executive coaching, emphasizing their accessibility and logical reasoning while highlighting the importance of human-AI collaboration to promote deep self-reflection. James et al. [james2024toward] evaluated LLMs for generating personalized goals in gamified mHealth apps. A study comparing LLM-generated and pre-determined goals highlighted the scalability and efficiency of LLM-driven content creation. Nepal et al. [nepal2024mindscape] introduced the MindScape app, which uses AI-generated journaling prompts tailored to smartphone behavioral data. The study showed improvements in mental well-being and self-reflection, emphasizing the app’s potential for enhancing traditional journaling for effective interventions that promote self-awareness and holistic well-being.
3 Methodology
@@
An AI chatbot was developed to generate EFT cues [ahmadi2024ai]; EFTeacher is modeled after the clinician-led EFT interviews [daniel2013future, hollis2020ecological, bickel2020does] and self-administered surveys [stein2021bleak, brown2023episodic, ruhi2022episodic] used previously. We implemented a website for a single-session study in which participants were given a username and password and could log into the website as shown in Figure 1 to complete the AI-assisted EFT interview. We used GPT-4-Turbo [achiam2023gpt] through Microsoft Azure, which is a university-approved cloud service, for EFTteacher. As shown in Figure 1, memory, in the form of a list, is the history of the conversation which is passed into the prompt along with the user query (message) so the chatbot can generate an output. As depicted in the diagram, the memory stores both the responses generated by the chatbot and the corresponding user queries. Each interaction between the human and GPT-4 is added to this list so that it can be used in the next input for the model.
@@
On the chat page of the website, depicted in Figure 2, users interacted with the chatbot, which guided them in generating a set of EFT cue texts. EFTeacher first introduced EFT cues to participants. They were then provided with clear criteria for crafting a good EFT cue. Participants were guided to ensure their cues met specific criteria: the event should be something they look forward to, vividly imaginable, confined to a single 24-hour period, and written in the present tense as if they are currently experiencing it.
Figure 2: Chatbot Interaction Page
To help them understand, participants were shown examples of well-crafted EFT statements that met all the criteria, as well as poor examples. They were then prompted to create their own EFT cues for specific timeframes, starting from one month into the future, ensuring their events were realistic, specific, and followed the given guidelines. This step-by-step approach ensured participants understood EFT cues and could create effective ones.
@@
For each time-frame, participants first provided brief contextual statements about the event. Then they answered a series of standardized questions to add specific details, such as where the event was taking place, who was present, what actions were occurring, and how they felt in that moment. EFTeacher then combined these brief statements and detailed responses into a coherent cue, typically 3-4 sentences long. If participants gave incomplete or unrelated answers, EFTeacher prompted them to revise their responses to ensure the quality and relevance of the cues. Once an EFT cue was generated, EFTeacher sought feedback from the participant to adapt the cue based on their preferences, ensuring it aligned with their personal perspective and needs. The user and chatbot interacted iteratively until a desired EFT cue was generated and accepted by the user. Once finalized, the EFT cue was extracted through a predefined routine and displayed in a dedicated box (cue box), as shown in Figure 2. As the conversation progressed and an EFT cue was generated for any time-frame, users could view the cues in the cue box, navigate back to previous cues, and review them. This allowed users to keep track of which events had already been discussed during the EFT generation dialog. For this study, once users completed covering events for seven different time-frames (one month, three months, one year, three years, five years, and ten years), the conversation concluded.
@@
Upon completing the conversation, the chatbot’s logs, including the entire conversation (referred to as memory), were saved as a file in the database. A designated routine stored the extracted EFT cues in the database, and then sent them to Qualtrics. There they evaluated the system’s usability using System Usability Scale (SUS) [brooke1996sus] questions. Also in Qualtrics, participants were presented with the EFT cue texts generated by EFTeacher, organized by the different time-frames. For each cue, participants were asked to evaluate its characteristics on a 5-point Likert scale, ranging from 1 (not at all) to 5 (extremely). They rated each cue based on four key aspects: Liking, Importance, Excitement, and Vividness. Afterward, the participants were presented with a survey to provide additional feedback about their experience by answering open-ended questions, as explained in detail in Section 3.2.